OpenStudy (anonymous):

Which is the simplified form of 16 times a to the eighth power all over 12 times a to the fourth power.? 4 over 3. a4 4a2 4a4 4 over 3. a2 please help Im doing a test and Ive got this question wrong about 4 to 5 times, please help.

OpenStudy (reemii):

\(\frac{16a^8}{12a^4}\) do you agree?

OpenStudy (anonymous):

yes,

OpenStudy (reemii):

simplify by separating \(\frac{16}{12}\times \frac{a^8}{a^4} = \frac43 a^4\).

OpenStudy (anonymous):

so when seperation do you divide the 16 and the 12 by 4?

OpenStudy (nurali):

4 over 3. a4

OpenStudy (reemii):

yes, 4 is a common factor of 16 and 12. the other fraction is done by the properties of the powers. \(a^k/a^m = a^{k-m}\).

OpenStudy (anonymous):

um ok but how exactly do you get that answer, i need to know how to do it please.

OpenStudy (anonymous):

sorry if that sounds kind of rude.

OpenStudy (reemii):

\(\frac{a^8}{a^4} = a^{8-4}\).

OpenStudy (anonymous):

oh ok reemii you've been a lot of help.

OpenStudy (reemii):

yw.

OpenStudy (anonymous):

ok while your here is it ok if i ask you another question, its one of the last ones i had a difficulty on,